OpenStudy (anonymous):

15(2 - x) = 13(3 - x)

OpenStudy (anonymous):

expand all parantheses by multiplying the terms inside the parentheses with the number in front of it 15(2 - x) = 30 -15 x do this for the entire equation and write result here

OpenStudy (anonymous):

X = - 9/2